# Implicit Subgoal Planning with Variational Autoencoders for Long-Horizon Sparse Reward Robotic Tasks

### v1

@@ VI-B Implementation Details
Figure 7: Subgoal sequences generated by option policy $\psi$ from the initial state to desired goal on AntMaze, PushOverObstacle, Stack, and Push4. The red sphere in AntMaze and transparent cubes in manipulation tasks are marked as subgoals of current timesteps.
The VAE-based Subgoal Generator is based on the CVAE structure and consists of an encoder $\psi$ and a decoder $\psi^{\prime}$. The encoder takes the current state and the desired goal as inputs and uses two probabilistic neural networks to output the mean $\mu$ and deviation $\sigma$. We then parameterize the Laplace distribution with the above outputs. We implement two probabilistic neural networks using one layer of fully connected network mapping from the encoder output to the subgoal space. The decoder has the opposite structure to the encoder. It takes as inputs the encoder-generated subgoals and the current state and outputs the reconstructed desired goal. The SAC is used as the underlying RL algorithm for action policy training and then fine-tuned online. We use a fixed temperature parameter $\alpha=0.01$ for all tasks. Table I presents the specifications of various parameters including the size of the replay buffer $\mathcal{D}$ and the mini-batch $\mathcal{B}$, the network structure, and the learning rates of SAC and VAESI. We set the number of subgoals used in Hindsight Sampler as Table I. The weight $\beta$ used in (13) refer to Table I. Both the option and action strategies use the Adam optimizer to update the network parameters. During the training process, we set a time limit of $T=30$ to update the subgoal when the current subgoal is not reached. We also relabel desired goals with 75% probability via future hindsight experience replay.
Bootstrapped data collection Deep reinforcement learning not only allows us to extract the best strategies from bootstrapped data but also allows robots to continually improve as they increasingly interact with spoiled data [43]. To accelerate training, we collect demo trajectories for offline policy training using a scripted policy that utilizes privileged information from the environment (e.g., object pose and fixture state). We collect $3000$ trajectories for offline policy training using scripted policies that utilize privileged information from the environment (e.g., object pose and fixture state). Within each demo trajectory, we initialize the environment randomly and set an arbitrary level of the desired goal. Trajectory lengths range from $100$ to $400$ for different tasks. The success rate of the scripted policies is no greater than $10\%$.
@@ VI-C Qualitative Analysis
Feasibility Fig. 7 shows the qualitative results of the subgoals generated by the four tasks. Inferred subgoals are marked as red spheres in Antmaze and as transparent cubes in the other manipulation tasks. Trajectories for the subgoals are generated based on the current state and the final goal, leading to the accomplishment of the final goal by resolving each subgoal. It should be noted that many of the inferred subgoals are noisy, indicating that certain subgoals may occasionally be unreachable within a given time step. This phenomenon is expected as these subgoals typically provide the direction to aim for, rather than the precise position the robot should achieve. The robot can ultimately achieve the final goal as this direction is progressively updated. Furthermore, the algorithm shows robustness towards any inaccurate subgoals that may emerge during strategy execution. For instance, the second subgoal generated in the PushOverObstacle task seems not feasible, but it indeed provides the direction for the robot to move. And with time goes on, when the time budget of the current option is over, the next subgoal will provide more accurate information for the robot to achieve.
@@
Figure 9: Evolution of $V$ values of subgoals in the PushOverObstacle task from the initial state to the desired goal. The last image integrates the entire evolution of the highest $V$ value in the subgoal space.
Optimality Utilizing the Value Selector module, we filter subgoals with higher $V$ values during task execution. To examine the effectiveness of the subgoals generated for the completion of long-horizon tasks, we randomly extract a sample of 1000 subgoals from the subgoal space of each task and compute their corresponding $V$ values. For the AntMaze environment, the entire area enclosed by the outer wall is identified as the subgoal space. The subgoal space for the other three manipulation tasks is designated as the aggregate of all positions of objects present on the table. We employ t-SNE[44], a tool that translates similarities between data points into joint probabilities, to reduce the high-dimensional data into a two-dimensional format. The results are then visualized in Fig. 8. If a subgoal has a lower $V$ value, it is marked with a lighter hue, indicating it may not be the optimal choice for the current state. In contrast, a subgoal with a higher $V$ value is shown with a darker hue, suggesting it would be more beneficial for the long-horizon task. Subgoal candidates obtained from the trained subgoal strategy are also plotted as green crosses in the figure. The plot reveals that the subgoals inferred by the VAE generator are mainly distributed within the space of higher $V$ values.
Additionally, we illustrate the evolution of the $V$ values of the subgoals in the PushOverObstacle task, transitioning from the initial state to the desired goal in Fig. 9. Each heatmap depicts the distribution of $V$ values in the subgoal space at the respective time step. The last heatmap presents the trajectory progression of the subgoals possessing the highest $V$ value within the subgoal space.
@@ VI-D Quantitative Analysis
Figure 10: The success rate of VAESI and other baselines. The performance of VAESI is higher than others in all four environments.
Figure 11: The episodic expected return during training. VAESI can achieve the highest expected return among all baseline methods.
TABLE II: Table of the expected return of different trained policies after testing 5 times with different seeds.
Policy AntMaze PushOverObstacle Stack Push4
RIS $-82.30\pm 2.57$ $-33.28\pm 10.93$ $-269.46\pm 20.34$ $-270.27\pm 9.81$
HP $-94.24\pm 2.39$ $-59.02\pm 6.19$ $-283.46\pm 11.26$ $-311.96\pm 27.90$
LEAP $-84.50\pm 2.85$ $-20.73\pm 10.43$ $-159.80\pm 26.69$ $-245.29\pm 35.65$
VAESI w/o HS $-93.25\pm 2.39$ $-99.04\pm 1.07$ $-284.67\pm 6.66$ $-344.58\pm 32.98$
VAESI w/o VS $-94.90\pm 2.20$ $-64.72\pm 4.84$ $-201.33\pm 29.71$ $-313.15\pm 35.63$
VAESI (Ours) -74.13 $\pm$ 2.74 -19.93 $\pm$ 8.34 -89.29 $\pm$ 16.44 -228.89 $\pm$ 11.97
@@
To ensure a fair comparison, all methods employ SAC as the underlying RL algorithm with identical network architecture. The trained policies are tested on four long-horizon tasks, inclusive of our method and the aforementioned baselines, with success rates depicted in Fig. 10. The results demonstrate that the VAESI algorithm, which utilizes both the Hindsight Sampler and the Value Selector module, yields the highest success rates. The strategy without the Hindsight Sampler, labeled VAESI w/o HS, encounters difficulties in rendering the subgoals generated by the RL agent feasible, resulting in low success rates across all four environments. Conversely, the strategy lacking the Value Selector module, denoted as VAESI w/o VS, still manages to achieve a measure of success (for instance, approximately a 74% success rate on the PushOverObstacle task). This suggests that the optimization module primarily functions as a facilitator to expedite the training of the subgoal generator, enabling the strategy to converge more rapidly and attain a higher success rate. When compared to other state-of-the-art subgoal-oriented reinforcement learning methods, VAESI consistently demonstrates a superior success rate. On the Push4 task, VAESI achieves a success rate exceeding 70%, while LEAP achieves a success rate of approximately 38%, and other methods yield success rates of less than 20%.
The evolution of the expected reward during training, i.e., the cumulative reward over the entire trajectory, is depicted in Fig. 11. After conducting five tests on four tasks using distinct seeds, the expected returns recorded for all the training policies are presented in Table II, with the highest return for each task emphasized in bold and the second highest return emphasized in underlining. A higher expected return indicates a superior policy. While all policies can potentially lead to higher expected returns, our approach produces the most substantial improvements in expected returns. For the AntMaze and PushOverObstacle tasks, all subgoal-oriented RL methods yield near-identical returns. However, in the case of the Stack3 and Push4 tasks, VAESI demonstrates significantly higher returns compared with other methods. The LEAP and HP methods rely solely on the initial state to infer subgoals, which proves insufficient for tasks with varying desired goals. These methods lack robustness for complex long-horizon tasks in most scenarios, and a change in the desired goal results in the generated subgoals becoming less accurate, leading to task failure. Furthermore, sampling subgoals randomly from the experience buffer can result in inaccurate subgoals, impeding the generation of optimal lower-level action policies.
@@ VI-E Real World Demonstrations
Figure 12: Environments setup with a Tiago++ robot arm, top-view camera, and blocks.
@@
Figure 13: The trajectories generated by the three manipulation tasks, PushOverObstacle, Stack3 and Push4, when executed in a real environment using the Tiago++ robotic arm. Here, green spherical shadows and transparent squares are used to represent the final goal and the subgoals generated during the execution.
@@
In order to demonstrate the adaptability of our algorithms in real-world, we employ the Tiago++ robotic arm [46] and its parallel fixtures to execute three long-horizon manipulation tasks: PushOverObstacle, Stack, and Push4. The setups of these three environments are the same as in the simulation. Our experiments utilize Pinocco [47] for both motion planning and inverse kinematics. As depicted in Fig. 12, red cubes are manipulated by the robotic arm to complete the assigned task, with the number of red blocks used in PushOverObstacle, Stack, and Push4 being 1, 3, and 4, respectively. Given the Tiago robot’s limited action and observation space, we mount a calibrated Intel RealSense D435i RGB-D camera on the top of the table to facilitate top-view observation. We also fix a soft beam to the end of the gripper to mitigate potential collisions between the gripper, the table, and the blocks during the execution of non-prehensile tasks, namely, PushOverObstacle and Push4.
@@
The strategy employed to guide the robot is identical to that used in the simulation. To gather environmental observations, we utilize Aruco markers for the detection of each object’s and gripper’s position. Fig. 13 illustrates the subgoals inferred from the initial state to the desired goals for the three manipulation tasks. The desired goals are denoted as green spherical shades and the subgoals as transparent squares.
During execution, the Aruco markers may obstructed by the robotic arm, resulting in the top-view camera being unable to obtain the position of markers at certain time steps, which is primarily exists in the Stack and Push4 task. To address this problem, we keep the position of the marker invariable and update it when markers can be detected again. Additionally, the robot arm is reset to a specific pose to observe the current state after the achievement of each subgoal. Thus, the top camera will eventually detect the uncertain markers with the gradual movement of the gripper, making it possible to complete the task.
Figure 14: The distance to desired goal from the current state when executing in the real world.
TABLE III: The success rate of three manipulation tasks in the real world.
Task PushOverObstacle Stack Push4
No interference 9/10 8/10 8/10
Interference after completion 10/10 8/10 10/10
Fig. 14 illustrates the evolution in distance between the current achieved goal and the final goal over the task execution. The task succeeds if the distance falls within a threshold of 0.05m. As the experiment goes on, the distance for all tasks steadily declines to 0.05. Given the greater complexity of Push4 in comparison to the other two tasks, it requires the most time to achieve success. We also conduct 10 trials for each task and tabulated the success rates in Table III. The success rate of three manipulation tasks refers to Table III.
Figure 15: The performance of the algorithm when the current state is externally disturbed by humans. The robot will update the subgoals in real time.
We also execute ten tests for each manipulation task, incorporating human disturbances to interfere with the completed task. The success rate of each task is detailed in Table III. It is shown that two non-prehensile tasks, PushOverObstacle and Push4, boast a 100% success rate, and the Stack task exhibits an 80% success rate. As depicted in Fig. 15, when a block is moved away from its designated goal by a human, the subgoal is promptly updated upon detection of the block, thereby guiding the gripper to complete the subtask. VAESI has the capacity to infer subgoals from arbitrary states because of the characteristics of sequential planning. Other subgoal generators, such as LEAP and HP, are not robust because they infer subgoals at the initial stage.
VI-F Failed Cases Analysis
Figure 16: The analysis of failed cases.
Although our algorithm generally performs well, occasional failures are observed, particularly in the Stack and Push4 tasks. We conduct 1000 tests on each of these tasks on different initial states and final goals, and quantify the failures across varying ranges of distance between the generated subgoal and the goal achieved at the end of the episodes. As shown in Fig. 16 (a), a notably high number of failures in the distance range of 00 to $0.1$ and above $1.0$, far surpassing the failure count in intermediate distance ranges. We analyze that the following two reasons may cause it. First, the subgoal is extremely close to the current state, as shown in Fig. 16 (b), indicating that the action policy prematurely deems the subgoal as achieved due to its minimal difference from the current state. Secondly, the subgoal is significantly far from the current state, as depicted in Fig. 16 (c), implying that the generated subgoal is still considered a long-horizon task, challenging the action policy to accomplish.
@@ VII Conclusion
In conclusion, we propose Variational Autoencoder-based Subgoal Inference (VAESI), a novel algorithm that leverages an explicit encoder model to produce feasible subgoals and an implicit decoder model for subgoal-generating acceleration, to solve the long-horizon task. The efficiency of the proposed algorithm is substantiated by both qualitative and quantitative results derived from simulations and real-world experiments, which highlight the superiority of our algorithm over prevailing baselines in terms of subgoal efficacy and final performance.

### v2

@@ VI-B Implementation Details
The Hybrid Subgoal Generator is based on the CVAE structure and consists of an encoder $\psi$ and a decoder $\psi^{\prime}$. The encoder takes the current state and the desired goal as inputs and uses two probabilistic neural networks to output the mean $\mu$ and deviation $\sigma$. We then parameterize the Laplace distribution with the above outputs. We implement two probabilistic neural networks using one layer of fully connected network mapping from the encoder output to the subgoal space. The decoder has the opposite structure to the encoder. It takes as inputs the encoder-generated subgoals and the current state and outputs the reconstructed desired goal. The SAC is used as the underlying RL algorithm for action policy training and then fine-tuned online. We use a fixed temperature parameter $\alpha=0.01$ for all tasks. Table I presents the specifications of various parameters including the size of the replay buffer $\mathcal{D}$ and the mini-batch $\mathcal{B}$, the network structure, and the learning rates of SAC and EISP. We set the number of subgoals used in Hindsight Sampler as Table I. The weight $\beta$ used in (11) refer to Table I. Both the option and action strategies use the Adam optimizer to update the network parameters. During the training process, we set a time limit of $T=30$ to update the subgoal when the current subgoal is not reached. For image-based observations, we employ a vector quantized variational autoencoder (VQ-VAE) [42] to extract features, which are subsequently used as input for subgoal generation. Detailed training procedures and results of the VQ-VAE can be found in Appendix -C and -D.
@@ VI-C Qualitative Analysis
Figure 7: Subgoal sequences generated by option policy $\psi$ from the initial state to desired goal on Stack, Push, OpenDrawer and Store. The transparent cubes in Stack and Push tasks, as well as the red and blue cubes in OpenDrawer and Store tasks, represent subgoals generated by EISP.
Figure 8: Subgoal sequences generated by option policy $\psi$ from the initial state to desired goal on image-based environments. The figures with yellow dashed rectangles (below) are the subgoals in the current observation (above).
Feasibility Fig. 7 shows the subgoals generated by EISP in four state-based tasks mentioned above. We mark the inferred subgoals as transparent cubes in the Stack and Push task and solid small cubes in the OpenDrawer and Store task. For each task, we show the sequence of subgoals in different stages, starting from the initial state to the final goal. Each task involves different skills; for example, the Stack task uses Pick and Place skills iteratively to reach the final goal, whereas the Store task requires four skills to finish the entire task. Besides, we also infer subgoals for image-based environments, specifically OpenDrawer and Store, with the results shown in Fig. 8. For these tasks, low-dimensional features extracted from image observations serve as input, and the inferred subgoals are reconstructed using a pretrained VQ-VAE decoder. These reconstructed subgoals are depicted in Figure 8 with yellow dashed rectangles.
It should be noted that many inferred subgoals are still noisy, indicating that certain subgoals may be unreachable within a given time step. This phenomenon is expected, as these subgoals typically provide the direction to aim for rather than the precise position the robot should achieve. The robot can ultimately achieve the final goal as this direction is progressively updated. Furthermore, EISP demonstrates robustness against non-optimal or unfeasible subgoals that may emerge during strategy execution, as it infers subgoals based on the current observation. For instance, the third subgoal generated in the Stack task seems not optimal since it is not directly on the path to the final goal. However, it still provides directional guidance for the robot. As time progresses and the time budget for the current option expires, the next subgoal will offer more accurate information for the robot to achieve its objective.
@@
Optimality To assess the optimality of generated subgoals, we visualize the distribution of subgoals in the high-dimensional subgoal space. For simplicity, the results of later experiments with OpenDrawer and Store were conducted in image-based environments. We compute the $V$ values of 1000 randomly sampled subgoals from the subgoal space of each task. The t-SNE [43] algorithm is then employed to reduce the high-dimensional data into a two-dimensional format. The results are visualized in Fig. 9. Subgoals with lower $V$ values are marked with lighter hues, indicating they may not be the optimal choice for the current state. In contrast, subgoals with higher $V$ values are shown with darker hues, suggesting they would be more beneficial for the long-horizon task. Subgoal candidates obtained from the trained subgoal strategy are also plotted as green crosses in the figure. The plot reveals that the subgoals inferred by the VAE generator are mainly distributed within the space of higher $V$ values. Initially, these subgoals are not located in regions with high $V$ values, but as the training goes on and the dataset is updated with high-quality data, they are gradually shifted to regions with high $V$ values.
Figure 10: The success rate of EISP and other baselines. The performance of EISP is higher than others in all four environments.
@@ VI-D Quantitative Analysis
To ensure a fair comparison, all methods employ SAC as the underlying RL algorithm with identical network architecture. The trained policies are tested on four long-horizon tasks, inclusive of our method and the aforementioned baselines, with success rates depicted in Fig. 10. The results demonstrate that the EISP algorithm, which utilizes both the Hindsight Sampler and the Value Selector module, yields the highest success rates. The strategy without the Hindsight Sampler, labeled EISP w/o HS, encounters difficulties in rendering the subgoals generated by the RL agent feasible, resulting in low success rates across all four environments. Conversely, the strategy lacking the Value Selector module, denoted as EISP w/o VS, still manages to achieve a measure of success (for instance, approximately a 60% success rate on the Stack task). This suggests that the optimization module primarily functions as a facilitator to expedite the training of the subgoal generator, enabling the strategy to converge more rapidly and attain a higher success rate. When compared to other state-of-the-art subgoal-oriented reinforcement learning methods, EISP consistently demonstrates a superior success rate. On the Push task, EISP achieves a success rate exceeding 70%, while LEAP achieves a success rate of approximately 38%, and other methods yield success rates of less than 20%.
TABLE II: Table of the expected returns for different trained policies after testing five times with different seeds. Higher returns indicate better performance.
Policy Stack Push OpenDrawer Store
RIS $-269.46$ $-270.27$256.374 $-865.453$
HP $-283.46$ $-311.96$ $-272.97$844.58
LEAP $-159.80$ $-245.29$ $-266.34$ $-639.97$
EISP w/o HS $-284.67$ $-344.58$ $-283.75$ $-990.43$
EISP w/o VS $-201.33$ $-313.15$ $-284.72$ $-775.61$
EISP (Ours) -89.29 -228.89 -246.91 -581.96
After conducting five tests on four tasks using distinct seeds, the expected returns recorded for all the training policies are presented in Table II, with the highest return for each task emphasized in bold and the second highest return emphasized in underlining. A higher expected return indicates a superior policy. While all policies can potentially lead to higher expected returns, our approach produces the most substantial improvements in expected returns. The LEAP and HP methods rely solely on the initial state to infer subgoals, which proves insufficient for tasks with varying desired goals. These methods lack robustness for complex long-horizon tasks in most scenarios, and a change in the desired goal results in the generated subgoals becoming less accurate, leading to task failure. Furthermore, sampling subgoals randomly from the experience buffer can result in inaccurate subgoals, impeding the generation of optimal lower-level action policies. For detailed information on episodic returns throughout the entire training procedure, you can refer to Appendix -E.
Figure 11: Environments setup. Left: Tiago robotic arm. Right: UR3 robotic arm.
@@ VI-E Real World Demonstrations
Figure 12: The trajectories generated by the four manipulation tasks when executed in a real environment are presented. For the Stack and Push tasks, green spherical shadows represent the final goals, and transparent squares denote the subgoals generated during execution.
@@
To demonstrate the adaptability of our algorithms in the real world, we employ the Tiago++ robotic arm [45] to execute Stack and Push tasks, and the UR3 robotic arm for OpenDrawer and Store tasks. The setups of these four environments are the same as in the simulation. Our experiments utilize Pinocchio [46] for both motion planning and inverse kinematics. As depicted in Fig. 11, red cubes are manipulated by the robotic arm to complete the assigned task, with the number of red blocks used in Stack and Push being 3 and 4, respectively. We mount a calibrated Intel RealSense D435i RGB-D camera on the top of the table to facilitate top-view observation. We also fix a soft beam to the end of the gripper to mitigate potential collisions between the gripper, the table, and the blocks during the execution of the Push task.
@@
The strategy employed to guide the robot is identical to that used in the simulation. To gather environmental observations, we utilize Aruco markers for the detection of each object’s and gripper’s position. Fig. 12 illustrates the subgoals inferred from the initial state to the desired goals for four manipulation tasks. The desired goals are denoted as green spherical shades and the subgoals as transparent squares. During execution, the Aruco markers may be occluded by the robotic arm, causing the top-view camera unable to detect the markers at certain time steps, primarily in the Stack and Push task. To address this problem, we maintain the last known position of the markers and update it when markers become detectable again. Additionally, the robot arm is reset to a specific pose to observe the current state after achieving each subgoal. We also conducted robustness tests and failed case studies, with the results presented in Appendix -F and -G.
@@ VII Conclusion
In conclusion, we propose Explicit-Implicit Subgoal Planning (EISP), a novel algorithm that leverages an explicit encoder model to produce feasible subgoals and an implicit decoder model that provides a guarantee on the worst-case log-likelihood of the subgoal distribution. The efficiency of the proposed algorithm is substantiated by both qualitative and quantitative results derived from simulations and real-world experiments, demonstrating its superiority over prevailing baselines in terms of subgoal efficacy and overall performance.